In redox reaction, H2O2 oxidizes K4[Fe(CN)6] into K+, Fe3+, CO3^2– and NO3– ions in acidic medium, than how many moles of H2O2 will react with 1 mole of K4[Fe(CN)6](a) 5 moles (b) 9 moles (c) 8 moles (d) 30.5 moles

Grade:12th pass

1 Answers

Shubhendu Dash
3 years ago
n factor of H2O2 = 2
give given n K4[Fe(CN)6] converted to K+,Fe+3,CO3-2,NO3factor
K+ remains K+ so n factor=0
Fe+2 changes to Fe+3 so n factor =1
C+2changes to C+4 so n factor =2
But as 6 carbon so n factor =6×2=12
N-3changes to N+5 so n factor=8
As there are 6 N so n factor =8×6=48
So the total n factor of K4[Fe(CN)6]=48+12+1+0=61
Total mole of H2O2 required to react with 1 mole of K4[Fe(CN)6] is 
=61/2=30.5